OpenStudy (anonymous):

I'm stuck on this question, and a little lost a bit in... Assume that p ≡ 3 (mod 4) and n ≡ x**2 (mod p). Given n and p, find one possible value of x. They give a hint to go with p = 3 + 4k, and to utilize Euler's Criterion to multiply by n at some point. But apart from some trivial statements, I'm stuck without introducing new variables I can't get rid of... Can anyone walk me through the first few steps of this?

OpenStudy (anonymous):

By Euler's Criterion, I can infer the following... n ** (p-1)/ 2 is congruent to 1 (mod p) x ** (p-1) is congruent to 1 (mod p)

OpenStudy (amistre64):

what does x**2 mean?

OpenStudy (anonymous):

Please excuse the notation. By x**2, I was attempting to show \[x ^{2}\]

OpenStudy (amistre64):

x^2, thats what i thougt but was trying to make sure

OpenStudy (amistre64):

p ≡ 3 (mod 4) p = 3 + 4k n ≡ x^2 (mod p) n = x^2 + pt n = x^2 + (3 + 4k)t n - (3 + 4k)t = x^2 for some given n and p

OpenStudy (amistre64):

not too sure if that would apply tho

OpenStudy (anonymous):

I can get this far, sadly. The introduction of the 't' variable drives me crazy, since now we don't know that...

OpenStudy (amistre64):

t and k are just some arbitrary integers that are used in the definition of "modity"

myininaya (myininaya):

Well \[x^2-a^2=(x-a)(x+a) => x^2=(x-a)(x+a)+a^2\] I this this could help. I used it to find an x.

myininaya (myininaya):

I translate for myself: "I think this could help."

OpenStudy (amistre64):

i cant say im familiar with Euler's Criterion to implement it in this case :/

myininaya (myininaya):

I don't think we need it.

OpenStudy (amistre64):

perhaps, i hate it when hints lead you astray ;)

myininaya (myininaya):

It is just a hint. We can ignore it or take it.

OpenStudy (amistre64):

i dunno about that ... last time i ignored a hint i ended up getting married :/

myininaya (myininaya):

I'm not familiar with euler's thingy ma-jigger either.

myininaya (myininaya):

I could be wrong actually.

myininaya (myininaya):

So I guess we are suppose to write x in terms of n and p.

OpenStudy (amistre64):

with any luck n>p

OpenStudy (amistre64):

n >= p that is

myininaya (myininaya):

Ok. Maybe we need the hint.

OpenStudy (anonymous):

Thanks for all the time you've given me. I'm not sure I'll ever find the answer the PSet wants me to, but I've mulled it over in my head for about two days now and reached the same conclusions that you have. You've restored my faith that I'm not in over my head as far as this material is concerned. Thanks!

myininaya (myininaya):

Is this from a book?

OpenStudy (anonymous):

It's from MIT OCW's Mathematics for Computer Science course. I could answer every question on the set but that with relative ease...

myininaya (myininaya):

So we can't choose values for p and n that satisfy the congruences? It has to have some kinda general solution for x.

OpenStudy (anonymous):

That's what it sounds like to me. I mean, I can satisfy it if I can choose n and p, according to those restrictions.

OpenStudy (anonymous):

The Problem Sets, unfortunately, don't have solutions, so I may never know what the intention of the question was -- but they're usually pretty good about wording them.

myininaya (myininaya):

If you ever find out the answer, I would like to see it.

OpenStudy (anonymous):

Well, currently the tiles in my bathroom are covered in white-board marker with every possible equation / congruence I can think of with this... so maybe once I give up, I'll take a shower and see it staring at me in the face.

myininaya (myininaya):

lol. That's interesting.

myininaya (myininaya):

I would have never guessed that.

OpenStudy (anonymous):

It's blocked here in China. Land of Piracy. Who'd have thought?

myininaya (myininaya):

This page actually shows how to get it. I'm trying to follow it.

OpenStudy (anonymous):

Holy moly! That page does show everything I need. I bow to your google-fu.

myininaya (myininaya):

lol. so apparently we have blum primes here which I never heard of such a cute name for primes. Do you understand it?

myininaya (myininaya):

How did they get (p+1)/4 as the exponent?

OpenStudy (anonymous):

Mulling that over as we speak. Is it because... n^((p-1)/2) * n = n^(p+1)/2 And n is the square of n?

OpenStudy (anonymous):

n is the square of x...

myininaya (myininaya):

So since \[n^\frac{p-1}{2} \underline{\approx } 1 \mod p \] and \[x^2 \underline{\approx} n (\mod p)\] then \[x^2 \underline{\approx} n ( 1) (\mod p)\] replacing this 1 here with \[x^2 \underline{\approx} n (n^\frac{p-1}{2} ) (\mod p)\] I guess this congruencey would still hold. \[x^2 \underline{\approx} (n^\frac{p+1}{2} ) (\mod p)\] Then somehow we just raise both sides to the 1/2 power to isolate x I guess.

myininaya (myininaya):

That would give us the desired result. But I wonder if we can do what I did above. Like I wonder if my logic is right.

OpenStudy (anonymous):

That's been my logic, as well...

myininaya (myininaya):

Your hint did say to multiply by that.

OpenStudy (anonymous):

I'm not sure how into Crypto you are, but I've found an interesting paper on the Blum-Blum-Shub Pseudo-Random Generator... http://www.cs.miami.edu/~burt/learning/Csc609.062/docs/bbs.pdf

myininaya (myininaya):

I do think crypto is interesting. I didn't get into as much of it as I wanted, well the math side anyways.

myininaya (myininaya):

well neither side really but i wanted the math part

myininaya (myininaya):

I think number theory is a cool subject, but I just have an elementary grasp on some of it.

OpenStudy (anonymous):

Yeah. I'm trying to "fix" the mistake of going into Liberal arts in my youth.